### v1

# Coupling quantum-like cognition with the neuronal networks within generalized probability theory

Abstract
The recent years are characterized by intensive applications of the methodology and mathematical apparatus of quantum theory, quantum-like modeling, in cognition, psychology, and decision making. In spite of the successful applications of this approach to a variety of psychological effects, e.g., the order, conjunction, disjunction, and response replicability effects, one may (but need not) feel dissatisfaction due to the absence of clear coupling to the neurophysiological processes in the brain. For the moment, this is just a phenomenological approach. In this paper we construct the quantum-like representation of the networks of communicating neurons. It is based not on standard quantum theory, but on generalized probability theory (GPT) with the emphasis of the operational measurement approach. We employ GPT’s version which is based on ordered linear state space (instead of complex Hilbert space). A network of communicating neurons is described as a weighted ordered graph that in turn is encoded by its weight matrix. The state space of weight matrices is embedded in GPT with effect-observables and state updates within measurement instruments theory. The latter plays the crucial role. This GPT based model shows the basic quantum-like effects, as e.g. the order, non-repeatability, and disjunction effects; the latter is also known as interference of decisions. This GPT coupling also supports quantum-like modeling in medical diagnostic for neurological diseases, as depression and epilepsy. Although the paper is concentrated on cognition and neuronal networks, the formalism and methodology can be straightforwardly applied to a variety of biological and social networks.
1Center for Mathematical Modeling in Physics and Cognitive Sciences
Linnaeus University, Växjö, SE-351 95, Sweden
2Center for Mathematical Science and Artificial Intelligence
Academy of Emerging Sciences, Chubu University, Kasugai, 487-8501, Japan
3Graduate School of Informatics, Nagoya University, Nagoya, 464-8601, Japan
4RIKEN Innovation Design Office, Hirosawa, Wako, 351-0198, Japan
5Felsenstein Medical Research Center, Petach Tikva, Israel
6Sackler Faculty of Medicine, Tel Aviv University, Tel Aviv, Israel
*Corresponding author email: Andrei.Khrennikov@lnu.se
keywords: quantum-like cognition, networks of communicating neurons, directed weighted graphs, generalized probability theory, order effect, interference effect
1 Introduction
The quantum information revolution (also known as the second quantum revolution) has not only the big technological impact (quantum computing, cryptography, simulators), but also the tremendous foundational consequences. One of them is understanding that the methodology and formalism of quantum theory can be applied outside of physics for modeling in cognition, psychology, decision making, biology including genetics and epigenetics, evolution theory, economics and finance, social and political sciences, game theory.111 See, for example, monographs [51, 20, 42, 9, 11, 43, 58, 12], the Palgrave handbook [44], and recent reviews [82, 57]. This area of research is known as quantum-like modeling (see, e.g., [81]).
It should be sharply distinguished from using of genuine quantum physics in cognition, as was done e.g. by Hameroff [40], Penrose [79], Vitiello [92]) and biology - “quantum biophysics” [6, 47]. The quantum-like models describe the behavior of macroscopic cognitive systems (e.g., human behavior) whose functioning can be modeled as information processing matching the laws of quantum information and probability and deviating from the classical laws.
During the last 10–15 years quantum-like studies flowered by covering new domains of science, for example, genetics and biological evolution theory [7, 8, 14]. Hence, the second quantum revolution was completed by the quantum-like revolution [81]. The main challenge for the quantum-like approach is that the neurophysiological mechanism of generation of the macroscopic quantum-like behavior is still unknown. The experimental statistical data collected in cognitive psychology is characterized by, e.g., the order, conjunction, disjunction, and response replicability effects, and contexuality. Such data can be successfully described by the quantum formalism. [^2] See, for example, papers [36, 19, 41, 94, 95, 53, 61, 17, 75, 76, 77, 24, 1, 2, 16, 31, 32, 23, 13, 30, 46, 33]. However, we still don’t have a proper mathematical model coupling the neurophysiological processes in the brain with the quantum-like description. In particular, this problem was highlighted in article [21]. Nowadays this is not just a theoretical concern. Recently the quantum-like methodology started to be used for the medical diagnostics of neurological diseases [85, 88] with the first “quantum-like patent” [86]. Although this method gives excellent predictions, the absence of the neurophysiological basis slows its application in hospitals, since physicians want to see its neurophysiological justification.333 In principle, one might refer to quantum biophysics and to genuine quantum physical processing in the brain, e.g., in the microtubules [40, 79]. The quantum-like approach doesn’t contradict the quantum-brain hypothesis. But this hypothesis has so many drawbacks and it was heavily criticized by many authors, e.g., by Tegmark [90]. From our viewpoint, its main drawback is that in the quantum-brain theory the neurons are not the basic units of mental information processing. To find such units, one should go to finer scales of space and time, up to the Planck scale in Penrose’s coupling of consciousness with quantum gravity. However, we want to stay firmly on the neuronal processing basis of cognition and consciousness.
In this paper we suggest a mathematical model of coupling between brain’s neurophysiology and the quantum-like project. It is based on the representation of functioning of brain’s neuronal networks within the generalized probability theory (GPT) (also known as the operational measurement/probability theory) [64, 38, 39, 25, 26, 67, 68, 72, 73], see [45, 80] for reviews. These networks are systems under analysis. Information exchange between neurons in a network $S$ is mathematically described by (generally directed) weighted graph $G_{S};$ neurons are mapped to the vertexes of $G_{S}$. We are well aware about difficulties of real experimental realization of such maps for the brain’s neuronal networks (see, e.g., [84]; so in our modeling we are still far from the experiment. Each graph $G_{S}$ corresponds to the information state of the network described by the weight matrix $\omega=(\omega_{ij})$. Matrix’s elements encode the strength of the signals between communicating neurons, e.g., the frequencies in the widely usedfrequency representation of the neural code [37] or correlations between the action The dynamics of weighted graphs mathematically describes the dynamics of network’s states. (But in this paper dynamics is not studied).
This is the good place to mention some previous works on the neural network background of the quantum-like models. De Barros and Suppes [28] (see also [29]) pointed out that neural oscillators can produce interference patterns like that obtained from unitary evolution. Takahashi and Cheon [89] proposed a nonlinear neural network that used squared amplitudes to compute the probability of response in a quantum model of cognition. Busemeyer et al. [21] outlined one way that a neural based system could perform the computations required by applications of quantum probability to human behavior. Recently quantum-like modeling on treelike graphs started to be used in the medical research [85, 88] reflecting the hierarchy in information processing by the brain. Moreover, one of the first works in quantum-like modeling of cognition [49] was on quantum mechanics of the treelike graphs which are mathematically encoded by $p$-adic numbers. Later $p$-adics was applied to quantum-like modeling in research on petroleum reservoirs [63](with the treelike structure of their capillary networks). We also mention article [55] in that the quantum-like representation of mental information is seeded to the distribution of action potentials of neurons interacting within networks. The present paper was influenced by the discussion with Gregory Scholes on quantum-like states of the neural networks described by the special class of graphs [83, 5]. We consider arbitrary graphs.
This paper is directed to those who are interested in the application of the quantum methods to cognition, psychology, and decision making. Generally researchers in this area are unaware about GPT. Moreover GPT is split in a few models that are apparently different but mathematically equivalent [67] (see [45, 80] for the reviews). We employ one special model; it is based on ordered linear spaces or so-called base norm spaces more precisely. To make the paper readable for experts in mathematical psychology, we briefly present the basics of GPT within the ordered linear space formalism in section 4.1.444It seems that the GPT approach can be more attractive for psychologists, than the standard formalism borrowed from quantum physics. GPT is “cleaned” from quantum mysteries and prejudices. This is simply a general mathematical formalism describing probabilities obtained in all kinds of observations done on all kinds of systems, physical, cognitive, social, financial. Before the formal introduction to GPT, we start by operating freely and as we hope clearly with the notions of GPT, and in section 3 we present the model bridging quantum-like cognition and brain’s networks of communicating neurons. Then we come back to this model in section 5 to show the order, interference (disjunction), and non-repeatability effects. For reader’s convenience the appendix starts with the GPT description of the conventional quantum probability (in complex Hilbert space).
This paper is conceptual, its aim is to present the concept of GPT coupling of neuronal activity with the quantum cognition project. In further papers we shall go deeper into the technical details. Although the paper is concentrated on cognition and neuronal networks, the formalism and methodology can be straightforwardly applied to a variety of biological and social networks with coupling to quantum-like modeling in genetics and social science [59, 4].
2 Graph representation of the networks of communicating neurons
We start with recollection [22] on the basics of brain’s neuronal signaling:
“The human brain is made up of approximately 86 billion neurons that talk to each other using a combination of electrical and chemical (electrochemical) signals. The places where neurons connect and communicate with each other are called synapses. Each neuron has anywhere between a few to hundreds of thousands of synaptic connections, and these connections can be with itself, neighboring neurons, or neurons in other regions of the brain. A synapse is made up of a presynaptic and postsynaptic terminal. The presynaptic terminal is at the end of an axon and is the place where the electrical signal (the action potential) is converted into a chemical signal (neurotransmitter release). The postsynaptic terminal membrane is less than 50 nanometers away and contains specialized receptors. The neurotransmitter rapidly (in microseconds) diffuses across the synaptic cleft and binds to specific receptors.”
The voltage-dependent channels endow the neuron with active signaling property. In particular, they enable a neuron to generate nerve impulses, also know as action potentials. In spite of variation in duration and amplitude as well as shape, they are typically treated as identical events. The duration of an action potential (approximately 1 ms) is ignored, an action potential sequence, or spike train, is represented by a series of all-or-none point events.
We note that each neuron has, on average, about 7,000 synaptic connections with other neurons. Axons vary in length, some of them are a millimeter or so, but the axons which go from the brain down the spinal cord, can extend for more than a meter. An axon has numerous side branches called axon collaterals, that is one neuron can send information to several others. The collaterals, as the roots of a tree, split into smaller extensions called terminal branches.
Consider some network $S$ in the brain consisting of $N$ neurons, e.g., a network performing some brain function or even the network of all neurons in the brain. Neuronal networks are systems of our quantum-like model. The symbol $S$ is used to denote a neuronal network of communicating neurons.
In the graph representation of $S$, neurons are mapped to the graph’s vertexes, connections between neurons are mapped to graph’s edges. We treat connections from the information viewpoint, so edges need not be physical edges consisting of axons and synapses. For example, edges (connections) between pairs of neurons are established by correlations between their signaling. We recall that if no two edges have the same endpoints we say there are no multiple edges, and if no edge has a single vertex as both endpoints we say there are no loops [91]. A graph $G_{S}$ representing network $S$ can have multiple edges and loops.
A directed graph consists of a set of vertices and a set of directed edges that are ordered pairs of vertices. A directed graph allows loops and allows two edges joining the same vertex, but going in the opposite direction. More than one edge going in the same direction between vertices, however, is not allowed [91]. The main model considered in this article (Model 1, section 3) is based on the directed neuronal graphs. Other two models are based on undirected graphs with loops, but without multiple edges.
Graphs describe the geometry of a network. But we model the “living community” of neurons communicating with each other. Therefore we consider the weighted graphs. If $G_{S}$ is the directed graph, then, for each pair of neurons $n_{i}$ and $n_{j}$, the weights $w_{ij}$ and $w_{ji}$ are assigned for edges $n_{i}\to n_{j}$ and $n_{j}\to n_{i}$. If $G_{S}$ is the undirected graph (without multiple edges), then the weight $w_{ij}=w_{ji}$ is assigned to any edge $(n_{i},n_{j})=(n_{j},n_{i})$.
2.1 Model 1: directed graph for neurons signaling
The weight $w_{ij}$ assigned to the graphs’ directed edge $n_{i}\to n_{j}$ can be interpreted as the probability that during some interval of time $\Delta$ (determining the scale of cognitive information processing in network $S)$, neuron $n_{i}$ sends a signal to neuron $n_{j}$. (This signal is of the electrochemical nature.) Denote this event $O_{i\to j}$. Thus, the neuronal network $S$ is represented by the weighted directed graph, each two neurons are connected by two arrow-edges, $n_{i}\to n_{j}$ and $n_{j}\to n_{i}$ arrows (two neurons cycle), and each neuron $n_{i}$ is feedback coupled to itself by the loop edge $n_{i}\to n_{i}$. No-signaling is encoded by the zero weight. So, we operate with a complete graph. This is convenient; we don’t appeal to the “hardware” of the connections between the neurons in the network $S$. By eliminating the edges with $w_{ij}=0$ we would obtain more interesting graph geometry, but this is the topic of further studies.
Such neural activity can be represented in the framework of classical probability theory; for example, as follows. Let $\xi_{ij}$ denote the classical random variable
$\xi_{ij}=\begin{cases}1,&\text{if}\;O_{i\to j}\;\text{holds},\\ 0,&\text{otherwise}.\end{cases}$ (1)
Then we set
$w_{ij}=E[\xi_{ij}]=P(\xi_{ij}=1).$ (2)
Here $0\leq w_{ij}\leq 1$. We note that generally $w_{ii}>0$ that is a neuron can send signals to itself. The GPT representation for this neuronal model will be presented in section 3.
2.2 Model 2: undirected graph for firing neural code
Instead of the model with the directed graph, we can consider the simpler model with undirected weighted graph. This model is connected with the neural code based on the firing frequency for neurons in the network $S$. For each neuron $n_{i}$ we define the random variable
$\xi_{i}=\begin{cases}1,&\text{if}\>n_{i}\;\text{sends spike},\\ 0,&\text{otherwise},\end{cases}$ (3)
that is in this model we are not interested in the redistribution of this spike between the neurons of the network $S$, as we do in the directed weighted graph model. Then we set
$w_{ij}=E[\xi_{i}\xi_{j}].$ (4)
The weight matrix $W_{S}=(w_{ij})$ is the covariance matrix of the random vector $(\xi_{1},...,\xi_{N})$.
2.3 Model 3: undirected graph for correlation between action potentials
The third model is also based on undirected graphs. For each neuron $n_{j}$ in a network $S$, its action potential is described by the real-valued random variable $\xi_{j}$. This is the continuous random variable representing neuron’s action potential as a physical potential and not as the discrete spike generation event. The weights of the graph $G_{S}$ are given by the pairwise correlations between the action (4). The weight matrix $W=(w_{ij})$ is the covariance matrix of the random vector $\xi=(\xi_{1},...,\xi_{N})$.
We shall construct a GPT representation for Model 1 (section 3). We postpone GPT design for Models 2 and 3 for a future publication. In the GPT terminology Model 1 has classical state space, but it shows the basic quantum-like effects, as the order, interference (disjunction), and non-repeatability effects.
3 GPT for directed neuronal graph (Model 1)
Denote the set of all $N\times N$ real matrices by the symbol $M\equiv M_{N}$. For matrix $W=(w_{ij})$, we define its norm
$||W||=\sum_{ij}|w_{ij}|.$ (5)
Hence, $M$ is normed space. Its dual space $M^{\star}$, the space of linear functionals, is algebraically isomorphic to $M$, so its elements are matrices $F=(f_{ij})$ with the dual paring $\langle F|W\rangle$ such that
$\langle F|W\rangle=\sum_{ij}f_{ij}w_{ij}$ (6)
for any $W=(w_{ij})\in M$. The norm on $M^{\star}$, is given by
$|||F|||=\max_{||W||=1}|\langle F|W\rangle|=\max_{ij}|f_{ij}|.$ (7)
For a matrix $W=(w_{ij})$, we write $W\geq 0$ if, for all $i,j$, $w_{ij}\geq 0$. We introduce a partial order structure, $W\geq W^{\prime}$ if $W-W^{\prime}\geq 0$. So, $M$ is normed ordered space. A linear functional $F$ is positive, $F\geq 0$, if $\langle F|W\rangle\geq 0$ for all $W\geq 0$, or equivalently $f_{ij}\geq 0$ for all $i,j$.
The directed weighted graph $G_{S}$ of neural network $S$ is represented by a matrix $W_{S}=(w_{ij}:w_{ij}\geq 0)$. For example, the matrix elements are given by probabilities (2). But, other neurophysiological entities can also be considered. To generate dimensionless entities, we normalize these graph-matrices and set
$\omega\equiv\omega_{S}=W_{S}/||W||.$ (8)
Thus, $||\omega_{S}||=1$. Denote the set of such matrices by the symbol $\Omega$ and this is the set of the states of the systems under consideration – the networks of communicating neurons:
$\Omega=\{\omega\in M:\omega\geq 0,||\omega||=1\}.$ (9)
We remark that the normalization condition implies that $\omega_{ij}\in[0,1]$. The set $\Omega$ is convex, closed, and bounded in $M$ (see Definition 1, section 4.2).
Now we turn to the issue of pure states, the extreme points of the convex set $\Omega$ (section 4.2). The set of pure states coincides with the set of matrices $(E_{ij})$, where $E_{ij}$ is the matrix with elements $e_{km;ij}=\delta(k-i)\delta(m-j)$. This is the basis in linear space $M$ consisting of the pure states. Any state $\omega$ can be represented uniquely as a convex combination (mixture) of pure states
$\omega=\sum_{ij}\omega_{ij}E_{ij}\,\;\mbox{where}\;\sum_{ij}\omega_{ij}=1,\omega_{ij}\geq 0.$ (10)
Thus, $\Omega$ is a simplex. The GPT-models based on such spaces are considered as classical [3, Theorem II.4.3]. In fact, classical-quantum interplay is more complex and the crucial role is played not by state spaces, but by the calculus of measurement instruments (section 4.3).
In GPT observables are defined as the collections of effects, positive linear functionals mapping state space into the segment $[0,1]$. Any linear functional $F=(F_{ij})$ acts on a mixture (10) as
$\langle F|\omega\rangle=\sum_{ij}F_{ij}\omega_{ij},$ (11)
so $F$ is an effect (i.e., $0\leq\langle F|\omega\rangle\leq 1$ for all $\omega$) if and only if $0\leq F_{ij}\leq 1$ for all $i,j$, or equivalently $F\geq 0$ and $|||F|||\leq 1$. As in the quantum formalism, for any $W\in M$,
$\langle F|W\rangle={\rm Tr}\;F^{\dagger}W,$ (12)
where $F^{\dagger}$ is the adjoint matrix of $F$. In GPT the important role is played by the unit effect $u$ such that $\langle u|W\rangle=||W||$ for $W\geq 0$; see (15) below). In the matrix model $u=(u_{ij}=1)$. This is the analog of the trace functional in quantum theory.
In general this is the basic scheme of GPT, states and effects. To proceed further, one should appeal to theory of measurement instruments. We want to do this properly and this needs more formal mathematical considerations that will be presented in the next section 4.
4 Generalized probability theory (operational measurement theory)
GPT was created in 1970th as an attempt to find the roots of the quantum formalism. Within GPT quantum formalism based on the complex Hilbert space is just a special version of the general operational formalism for all kinds of measurements. It unifies the classical and quantum measurement theories and opens the door to the variety of quantum-like formalisms. It is clear that probabilistically non-classical properties of mental phenomena need not imply that they should be described by the quantum (complex Hilbert space) formalism, quantum probability, information and measurement theory. Some GPT-based models can match better cognitive information processing and decision making; this was highlighted in [60]. [^5] In 2015 one of the coauthors (AKH) by staying in Vienna had long discussions with Anton Zelinger on possible applications of quantum theory to cognition. These discussion were initiated by conversations about his debates with Dalai Lama on spiritual versus quantum mechanical phenomena. He was strongly in favor of appealing to generalized theories which are neither classical nor quantum; see also his preface to book [81]. He also guessed that it may be that final quantum-like theory of cognition would be mathematically so complicated and exotic that the complex Hilbert formalism would not more look as complex and advanced. But this is not the case for the model suggested in this article which is “less quantum” than the quantum mechanics.
In this paper we employ GPT based on theory of ordered linear spaces [25, 26, 67, 68, 72, 73, 74]. In contrast to the majority of GPT-studies, we emphasize not the structures of the state and effect spaces, but the calculus of measurement instruments giving the probabilities (via generalized Born’s rule) and states’ updates generated by the measurements feedback. This approach to GPT started by Davies and Lewis [25] and known as quantum instruments theory, and it has been developed by Ozawa [67, 68, 69, 70, 71, 72, 73, 74], see also the recent article by Khrennikov [60] in that this approach was restructured in the contextual framework. Such contextual viewpoint of GPT and quantum instruments theory enlightens the possibility of their applications to cognitive psychology, decision making, social science [60]. GPT research which is concentrated on generalization of the notions of state and observable was started in [64] and essentially developed by Gudder [38, 39]. Ozawa [67]. established the equivalence between Gudder’s and Davies–Lewis’ approaches.
4.1 Ordered linear spaces
For simplicity, we restrict considerations to finite dimensional linear spaces. Let $X$ be a real linear space. A subset $C$ of $X$ is called a cone, if for any positive real number $r$, $rC\subset C$. A set $B$ is a base of the cone $C$ if, for every $x\in C$, there is a unique $\lambda>0$ such that $\lambda x\in B$ or in other words all elements of $C$ can be obtained via scaling of elements of $B$ with positive scalars.
Let $X$ be a real linear space with a partial order $\leq$ that is consistent with the linear space structure: for any $x,y,z\in X$ and real $r\geq 0,$ if $x\leq y$, then $x+z\leq y+z$ and $rx\leq ry$. Such $X$ is called an ordered linear space. Set $X_{+}=\{x\in X:x\geq 0\}$. This set is a convex cone containing zero. The elements of $X^{+}$ are called positive. If $x,y\in X$, then $x\leq y$ if and only if $y-x\in X^{+}$.
Consider the dual space $X^{\star}$, the space of linear functionls on $X$. Since we work in the finite dimensional case, spaces $X$ and $X^{\star}$ are algebraically isomorphic. A linear functional is called positive if $f(x)\equiv\langle f|x\rangle\geq 0$ for any $x\geq 0$. The set of all positive functionals $X^{\star}_{+}$ is a convex cone in $X^{\star}$. It is called the dual cone of $X_{+}$. The order structure determined by $X^{\star}_{+}$ coincides with the point wise order structure on $X^{\star}$, i.e., $f\leq g$ iff $f(x)\leq g(x)$ for all $x\in X_{+}$. To use analysis, a linear space should be endowed with a topology, e.g., determined by a norm |||||$\cdot$||| | ⋅ | | matching the order structure. Let $X$ be an ordered linear space with norm |||||$\cdot$||| | ⋅ | | such that there exists a real number $k>0$ such that, for all $x,y\in X,0\leq x\leq y$ implies that
$||x||\leq k||y||.$ (13)
Such $X$ is called a normed ordered space. The matrix space $M$ with norm (5) and its dual space $M^{\star}$ with norm (7) are normed ordered space with $k=1$.
In any normed space the symbols $B_{r}(a)$ and $S_{r}(a)$ denote the ball and sphere of radius $r>0$ centered at the point $a$. The dual space is endowed with the norm $||f||_{\star}=\max_{||x||=1}|\langle f|x\rangle|$. The corresponding balls and spheres are denoted as $B^{\star}_{r}(a)$ and $S^{\star}_{r}(a)$.
4.2 States and observables
Definition 1. [80] State space $\Omega$ is a convex, closed, and bounded subset of a real, finite-dimensional vector space with Euclidean topology.
The above definition is presented just to simplify the mathematical formalism. We note that such set $\Omega$ is compact. We adjust this definition to the ordered space approach. Let $X$ be an ordered normed space. A positive linear functional $f$ on $X$ is called strictly positive if $f(x)>0$ for all $x\in X_{+}\setminus\{0\}$. A normed ordered space $X$ is called a base norm space if $X_{+}$ is closed and there exists a strictly positive linear functional $u$ such that $u(x)=||x||$ for all $x\in X_{+}$ [66, 67]. In this case $X_{+}\cap S_{1}(0)=\{x\in X_{+}\mid||x||=1\}$ is called the base of $X$, which is a convex closed, and bounded base of the cone $X_{+}$ of positive elements. The matrix space $M$ with norm (5) is a base norm space but its dual space $M^{\star}$ with norm (7) is not.
Definition 1a. State space $\Omega$ is the base of a finite-dimensional base norm space $X$, i.e., $\Omega=X_{+}\cap S_{1}(0)$.
Let $\Omega$ be a state space and let $x\in\Omega$. It is said that $x$ is an extreme point of $\Omega$, or equivalently that $x$ is a pure state, if for every $y,z\in\Omega$ and $\lambda\in(0;1)$ such that $x=\lambda y+(1-\lambda)z$ we have $x=y=z$. And a state space $\Omega$ is a polytope if it has finitely many pure states.
A positive functional, $E\in X^{\star}_{+}$, is called an effect on the state space $\Omega$ if
$0\leq\langle E|\omega\rangle\leq 1\;\mbox{for \;all}\;\omega\in\Omega.$ (14)
Effects are the basic observables in the operational theories of measurement. They can be described solely in terms of functions on the state space $\Omega$, as linear space extensions of affine functionals defined on $\Omega$ and valued in the segment $[0,1]:$
$\langle f|\lambda\omega_{1}+(1-\lambda)\omega_{2}\rangle=\lambda\langle f|\omega_{1}\rangle+(1-\lambda)\langle f|\omega_{2}\rangle$
for $\lambda\in[0,1]$ and $\omega_{1},\omega_{2}\in\Omega$. The set of effects is denoted by the symbol ${\mathcal{E}}$. This set is convex and compact. It follows from Definition 1a that ${\mathcal{E}}$ contains the unit effect $u$, for $x\in X_{+}$,
$\langle u|x\rangle=1\;\mbox{if and only if}\;x\;\mbox{belongs to}\;\Omega,$ (15)
that is $\langle u|x\rangle=||x||$ for all $x\in X_{+}$ and
$\Omega=\{x:x\geq 0,\langle u|x\rangle=1\}\;$ (16)
In terms of the unit functional, states can be characterized as positive vectors constrained by equality (15).
Let $(E_{j})$ be a sequence of effects such that
$\sum_{j}E_{j}=u,\;\text{that is}\;\sum_{j}\langle E_{j}|\omega\rangle=\langle u|\omega\rangle=1,$ (17)
and let $(\alpha_{j})$ be a sequence of real numbers. An observable $A$ is a sequence of couples $(\alpha_{j},E_{j})$, where set $V_{A}=(\alpha_{j})$ encodes the outcomes of observations (“spectrum” of the observable) and the effects determine the probabilities of these outcomes via the generalized Born’s rule:
$P(A=\alpha_{j}|\omega)=\langle E_{j}|\omega\rangle,\;\omega\in\Omega.$ (18)
The average outcome of observable $A$ in the state $\omega$ is as usual via the formula:
$\langle A\rangle_{\omega}=\sum_{j}\alpha_{j}P(A=\alpha_{j}|\omega)=\langle\sum_{j}\alpha_{j}E_{j}|\omega\rangle.$ (19)
Thus, formally the observable $A$ can be described as the functional $A=\sum_{j}\alpha_{j}E_{j}\in X^{\star};$ then $\langle A\rangle_{\omega}=\langle A|\omega\rangle$. But this description is useful only for calculation of averages
4.3 Instruments
Let $J:X\to X$ be a linear operator preserving the positive cone, i.e., $J(X_{+})\subset X_{+}$. Its dual operator $J^{\star}:X^{\star}\to X^{\star}$ also preserves positivity: let $x\in X_{+},f\in X^{\star}_{+}$, then $\langle J^{\star}f|x\rangle=\langle f|Jx\rangle\geq 0$, since $Jx\geq 0$.
Let $A=(\alpha_{k},E_{k})$ be an observable with the discrete set of outcomes $V_{A}=(\alpha_{k})$ and the family of effects $(E_{k})$ and let, for each outcome $\alpha\in V_{A}$, there is defined a linear operator $J(\alpha):X\to X$ preserving the positive cone $X_{+}$. It is assumed that the family of maps $(J(\alpha))$ is constrained as
$J(V_{A})\equiv\sum_{\alpha}J(\alpha):\Omega\to\Omega.$ (20)
We remark that this constrain implies that
$\langle u|J(V_{A})x\rangle=\langle u|x\rangle,\;x\in X_{+}.$ (21)
For $x\in X_{+}$, set $\omega_{x}=x/\langle u|x\rangle$. Since $\langle u|\omega_{x}\rangle=1$, we have that $\omega_{x}\in\Omega$ and, hence, $\langle u|J(V_{A})\omega_{x}\rangle=1$, and (21) holds.
Hence,
$0\leq\langle u|J(\alpha)\omega\rangle\leq 1$ (22)
We remark that the last part of equality (21) implies that
$\langle u|J(V_{A})\omega\rangle=1.$ (23)
In operational measurement theory, measurements with the outcome $\alpha$ induce the state-update transform of the state space $\Omega:$
$\omega\to\omega_{\{A=\alpha\}}=\frac{J(x)\omega}{\langle u|J(\alpha)\omega\rangle}.$ (24)
We remark that $\langle u|\omega_{\{A=\alpha\}}\rangle=1$, hence $\omega_{\{A=\alpha\}}\in\Omega$, see (15).
For each $J(\alpha):X\to X$ consider the dual operator $J^{\star}(\alpha):X^{\star}\to X^{\star};$ it is positivity preserving. Hence, $J^{\star}(\alpha)u\in X^{\star}_{+}$. Moreover, $\langle J^{\star}(\alpha)u|\omega\rangle=\langle u|J(\alpha)\omega\rangle$. Hence,
$\sum_{\alpha}\langle J^{\star}(\alpha)u|\omega\rangle=\langle u|\sum_{x}J(\alpha)\omega\rangle=\langle u|J(V_{A})\omega\rangle=1.$ (25)
[…]
$E(\alpha)=J^{\star}(\alpha)u$ (26)
is an effect. Moreover, equality (25) implies that, for any $\omega\in\Omega$,
$\langle\sum_{\alpha}E(\alpha)|\omega\rangle=\sum_{\alpha}\langle J^{\star}(\alpha)u|\omega\rangle=1.$ (27)
Since $u$ is uniquely characterized by the property, $\langle u|\omega\rangle=1$ for any $\omega\in\Omega$, equality (27) implies that $\sum_{\alpha}E(\alpha)=u$, i.e., the system $A=(\alpha,E(\alpha))$ is an observable.
Definition 2. A family of positivity preserving linear operators $J(\alpha):X\to X,\;\alpha\in V_{A}$, constrained as (20) is called a measurement instrument.
Each instrument generates an observable $A=(\alpha,E(\alpha))$ with the effects given by the equality (26); the probabilities of outcomes are determined by the generalized Born’s rule:
$P(A=\alpha|\omega)=\langle E(\alpha)|\omega\rangle=\langle u|J(\alpha)\omega\rangle,$ (28)
measurement’s back action on a state $\omega$ is determined by formula (24).
Definition 3. Generalized probability theory (GPT), or operational measurement theory, is a triple $(X,\Omega,\mathfrak{J})$, where $X$ is a base norm space, $\Omega$ is a state space, and $\mathfrak{J}$ is the set of measurement instruments generating observables (families of effects) by the rule (26).
As was emphasized in [76], the basis of any measurement model is its instrument calculus. One can’t start from observables, as often done, in quantum information theory, where one is typically fine by introducing just observables. An observable is only a byproduct, or an “effect”, of an instrument, and the statistics of a measurement is fully described by the corresponding instrument.
4.4 Conditional and sequential joint probability distributions
Within GPT, operational measurement theory, we can define conditional probability [60, 69, 70]. Let $(J_{A}(\alpha))$ and $(J_{B}(\beta))$ be two instruments generating the observables $A$ and $B$. The probability to obtain the output $A=\alpha$ under the condition that the output $B=\beta$ was obtained in the preceding measurement of observable $B$ is given by the formula:
$P(A=\alpha|B=\beta,\omega)$ $=$ $P(A=\alpha|\omega_{\{B=\beta\}})=\langle E_{A}(\alpha)|\omega_{\{B=\beta\}}\rangle$ (29)
$=$ $\langle u|J_{A}(\alpha)\omega_{\{B=\beta\}}\rangle=\frac{\langle u|J_{A}(\alpha)J_{B}(\beta)\omega\rangle}{\langle u|J_{B}(\beta)\omega\rangle}$
$=$ $\frac{\langle u|J_{A}(\alpha)J_{B}(\beta)\omega\rangle}{P(B=\beta|\omega)}.$
By using conditional probability we define the sequential joint probability
$P_{BA}(\beta,\alpha|\omega)=P(B=\beta|\omega)P(A=\alpha|B=\beta,\omega)=\langle u|J_{A}(\alpha)J_{B}(\beta)\omega\rangle.$ (30)
Sequential measurement of observables $A_{1},...,A_{n}$ corresponding to instruments $J_{A_{j}}$ generates the joint probability distribution $P_{A_{1}...A_{n}}(\alpha_{1},...,\alpha_{n})=\langle u|J_{A_{n}}(\alpha_{n})...J_{A_{1}}(\alpha_{1})\omega\rangle.$
4.5 Order effect
The question order effect is well studied in psychology and sociology, e.g., for social opinion pools [65]. In classical probability theory the order effect is absent. In [94, 95] it was represented in the quantum-like framework based on the projection type (von Neumann [93]) state updates. But then in [53] it was shown that such projection description can’t be combined with another psychological effect, the response replicability effect. To resolve this problem, in [75, 76] the calculus of quantum instruments was employed. We now consider the order effect within GPT.
Consider two observables (in psychology - questions, tasks) $A,B$ associated with the instruments $J_{A},J_{B}$. They show the order effect in a state $\omega$, if for some outcomes $\alpha_{1}$ and $\alpha_{2}$,
$P_{AB}(\alpha,\beta|\omega)\not=P_{BA}(\beta,\alpha|\omega),$ (31)
that is
$\langle u|(J_{B}(\beta)J_{A}(\alpha)-J_{A}(\alpha)J_{B}(\beta))\omega\rangle\not=0,$
that is the instrument operators don’t commute in this state (cf. [76]),
$[J_{A}(\alpha),J_{B}(\beta)]\omega\not=0.$ (32)
4.6 Interference effect
We remark that the interference of probabilities is one of the basic quantum effects; for example, this effect can be observed in the two slit experiment (as was emphasized by Feynman [34, 35], see also [48]). In cognitive science and decision making it was analyzed in [50] as a nonclassical property of cognitive information processing, its coupling with the disjunction effect in cognitive psychology was established in [19]. We remark that in classical probability theory probabilities don’t interfere.
As was explained in [48], quantum-like interference of probabilities can be written as the violation of one of the basic laws of classical probability theory, the formula of total probability (FTP). For dichotomous observables $A=\alpha_{1},\alpha_{2}$ and $B=\beta_{1},\beta_{2}$, FTP has the form: $P(A=\alpha)=\sum_{\beta}P(B=\beta)P(A=\alpha|B=\beta).$ In classical probability the role of the state is played by the probability measure of the Kolmogorov probability space, to match our formalism denote this Kolmogorovian probability by the symbol $\omega$, then FTP is written as
$P(A=\alpha|\omega)=\sum_{\beta}P(B=\beta|\omega)P(A=\alpha|B=\beta,\omega).$ (33)
In an operational measurement model the left-hand side of FTP equals $P(A=\alpha|\omega)=\langle u|J_{A}(\alpha)\omega\rangle$ and the right-hand side of FTP equals
$\sum_{\beta}P(B=\beta|\omega)P(A=\alpha|B=\beta,\omega)$ (34)
$=$ $\sum_{\beta}P_{BA}(\beta,\alpha|\omega)=\sum_{\beta}\langle u|J_{A}(\alpha)J_{B}(\beta)\omega\rangle$
$=$ $\langle u|J_{A}(\alpha)J_{B}(V_{B})\omega\rangle.$
The interference term is defined as the difference between these two expressions
$\delta_{BA,\omega}(\alpha)=\langle u|J_{A}(\alpha)(I-J_{B}(V_{B}))\omega\rangle.$ (35)
If $\delta_{BA,\omega}(\alpha)=0$ we have classical FTP (for this $\beta)$, if $\delta_{BA,\omega}(\alpha)>0$, this is constructive interference, it increases the probability (this is the basic tool of increasing effectiveness of quantum computing comparing with classical one), if $\delta_{BA,\omega}(\alpha)<0$, then interference is destructive.
Suppose that
$[J_{A}(\alpha),J_{B}(V_{B})]\omega=0,$ (36)
in this case $\langle u|J_{A}(\alpha)J_{B}(V_{B})\omega\rangle=\langle u|J_{B}(V_{B})J_{A}(\alpha)\omega\rangle.$ Since $J_{A}(\alpha)\omega\geq 0$, by equality (21) $\langle u|J_{B}(V_{B})J_{A}(\alpha)\omega\rangle=\langle u|J_{A}(\alpha)\omega\rangle$. Therefore instrument commutativity condition (36) implies that the interference term is zero, $\delta_{BA,\omega}(\alpha)=0$.
4.7 Repeatable measurements
Consider observable $A=(\alpha,E_{\alpha})$ based on instrument $J$ such that for some $a$, $J(a)$ is a positive projection that is $J(a)\geq 0$ and $J(a)^{2}=J(a)$. Then, for any state $\omega$, $P(A=a|A=a,\omega)=P(A=a|\omega_{\{A=a\}})=\langle u|J(a)\omega_{\{A=a\}}\rangle=\langle u|J(a)^{2}\omega\rangle/\langle u|J(a)\omega\rangle=1$. We also have $J(V_{A})=\sum_{\alpha}J(\alpha):\Omega\to\Omega$. Thus, $\sum_{\alpha}P(A=\alpha|A=a,\omega)=\langle u|J(V_{A})\omega_{\{A=a\}}\rangle=1$. Consequently,
$P(A=\alpha|A=a,\omega)=\delta_{\alpha,a}$ (37)
Hence, for the $a$-outcome the measurement is repeatable. However, if, for some $a,J(a)$ is not projection, i.e., $J(a)^{2}\not=J(a)$, then generally $P(A=a|A=a,\omega)=\langle u|J(a)^{2}\omega\rangle/\langle u|J(a)\omega\rangle\not=1$.
We remark that in classical probability theory all measurements (they are given by random variables) are repeatable; by the Bayes formula for conditional probability, $P(A=\alpha|A=a,\omega)=P(A=\alpha,A=a|\omega)/P(A=a|\omega)=\delta_{\alpha,a}$. We also remark that in quantum probability theory the projection type measurements are repeatable.
5 Model 1: Order, interference, and repeatability effects
In matrix space $M_{2}$ consider observables $A$ and $B$ with the values $\alpha_{1},\alpha_{2}$ and $\beta_{1},\beta_{2}$ with instruments realized as multiplication by matrices:
$J_{A}(\alpha_{1})=\begin{vmatrix}1&0\\ 0&0\\ \end{vmatrix},J_{A}(\alpha_{2})=\begin{vmatrix}0&0\\ 0&1\\ \end{vmatrix};J_{B}(\beta_{1})=\begin{vmatrix}0&1\\ 0&0\\ \end{vmatrix},J_{B}(\beta_{2})=\begin{vmatrix}0&0\\ 1&0\\ \end{vmatrix},$
then $J_{A}(V_{A})=\begin{vmatrix}1&0\\ 0&1\\ \end{vmatrix},\;J_{B}(V_{B})=\begin{vmatrix}0&1\\ 1&0\\ \end{vmatrix}$ and, for $\alpha_{1}$ and $\beta_{1}$, we have $J(\alpha_{1})J(\beta_{1})=J(\beta_{1}),J(\beta_{1})J(\alpha_{1})=0$ that is $[J(\alpha_{1}),J(\beta_{1})]=J(\beta_{1})\not=0$.
For an arbitrary state $\omega,\;P_{BA}(\beta_{1},\alpha_{1})=\langle u|J(\alpha_{1})J(\beta_{1})\omega\rangle=\langle u|J(\beta_{1})\omega\rangle=\omega_{21}+\omega_{22}\not=P_{AB}(\alpha_{1},\beta_{1})=0$. Thus, the order effect can be found for any network of two neurons $(n_{1},n_{2})$ in that $n_{2}$ is active.
[…]
Now turn to the problem of repeatability of observations. We note that $J(\beta_{1})^{2}=0$, i.e., it is not projection; then $P(B=\beta_{1}|B=\beta_{1},\omega)=0$. Therefore $B$ measurement is not repeatable. In contrast the $A$-measurement is repeatable.
6 Model 1: Compound systems, tensor product representation
Consider now two neuronal networks $S_{k},k=1,2$, containing $N_{k}$ neurons respectively, and consider the compound system $S=(S_{1},S_{2})$. In quantum theory and GPT this system should be described by the tensor product construction based on the linear space $M\equiv M_{N_{1}N_{2}}=M_{N_{1}}\otimes M_{N_{2}}$. How can one couple this tensor product construction to our neuronal networks model? Straightforwardly one treats $S$ as the network with $N_{A}+N_{B}$ communicating neurons and represents $S$ by matrix of the size $(N_{1}+N_{1})^{2}$, and not of the size $(N_{1}N_{2})^{2}$ as GPT requieres. The essence of our further consideration is in understanding that the state of a compound system should be interpreted differently from the state of a single system. This is one of the consequences the studies on so called prequantum classical statistical field theory [52], an attempt to model quantum phenomena with classical random fields. In this article we proceed with this ideology which matches well with our neuronal model.
We turn to the concrete representation of the weights $w_{ij}$ via averages of the classical random variables $\xi_{ij}$, see (1), (2). Denote these variables for networks $S_{1},S_{2}$ as $\xi^{(1)}_{ij},\xi^{(2)}_{ij}$. Consider the covariance matrix for these variables
$w_{ijkm}=E[\xi^{(1)}_{ij}\xi^{(2)}_{km}].$ (38)
The matrix $W_{S}=(w_{ijkm})$ belongs to the matrix space $M\equiv M_{N_{1}N_{2}}=M_{N_{1}}\otimes M_{N_{2}}$. In this case, $w_{ijkm}$ represents the joint probability of the event that both $O^{(1)}_{i\to j}$ and $O^{(2)}_{k\to m}$ occur simultaneously, i.e.,
$w_{ijkm}=P(\xi^{(1)}_{ij}=1,\xi^{(2)}_{km}=1).$ (39)
We remark that the basis of linear space $M$ can be selected as $E_{ijkm}=E_{ij}\otimes E_{km}$. Each $W\in M$, can be represented as the linear combination $W=\sum_{ijkm}w_{ijkm}E_{ijkm}$. The norm on $M$ is defined as $||W||=\sum_{ijkm}|w_{ijkm}|$. A matrix $W$ is positive if all its elements are non-negative, i.e., the positive cone $C_{+}=\{W:w_{ijkm}\geq 0\}$. So, matrices generated by the compound neuronal network via (38) are positive. The state space of the compound system $S=(S_{1},S_{2})$ is defined as in (9), $\Omega=\{\omega:\omega\geq 0,||\omega||=1\}$. The matrix $W_{S}$ of the network $S$, see (38), is transformed to the corresponding state via normalization, $W_{S}\to\omega=W_{S}/||W_{S}||$.
Thus, the GPT description of the compound network via the covariance matrix differs from the GPT description of the network $S_{1}\cup S_{2}$ composed of all neurons belonging to $S_{1}$ and $S_{2}$ and connected by the directed edges . For $S_{1}\cup S_{2}$ description, we need to establish all connections between the neurons $n_{i}^{(1)}\in S_{1}$ and $n_{j}^{(2)}\in S_{2}$ and assign to the corresponding directed edges the weights. The $S_{1}\otimes S_{2}$ description doesn’t contain these connections between neurons of $S_{1}$ and $S_{2}$. Only the cross-correlations between signals inside these networks are taken into account. In this sense the $S_{1}\otimes S_{2}$ description is incomplete, since it doesn’t recount connections between the neurons of the networks $S_{1}$ and $S_{2}$. On the other hand, it recounts correlations, or more specifically the concurrences, between signals within the network $S_{1}$ and the network $S_{2}$.
The graph of the $S_{1}\otimes S_{2}$ description can be constructed as follows. Denote the edges of the directed graphs $G_{S_{p}},p=1,2$, by the symbols $e^{(p)}_{ij}$. These are the vertexes of the graph under construction that is symbolically denoted as $G_{S_{1}\otimes S_{2}}$. This graph is undirected and weighted, and any vertex $e^{(1)}_{ij}$ is connected with any vertex $e^{(2)}_{km}$ by the edge $e^{(12)}_{ijkm}$ with the weight $w_{ijkm}$. For example, for $N_{1}=2,N_{2}=3$, the graph $G_{S_{1}\otimes S_{2}}$ has $32$ (undirected) edges and the graph $G_{S_{1}\cup S_{2}}$ has 25 (directed) edges. The graph $G_{S_{1}\otimes S_{2}}$ is not the standard tensor product $G_{S_{1}}\otimes G_{S_{2}}$ of the graphs $G_{S_{1}}$ and $G_{S_{2}}$. The latter graph will be used for the description of compound systems in models 2, 3 (see appendix). The graph $G_{S_{1}\otimes S_{2}}$ can be called the edge tensor product of the directed graphs $G_{S_{1}}$ and $G_{S_{1}}$.
Suppose now that signals in $S_{1}$ and signals in $S_{2}$ are independent, i.e., the corresponding random variables are independent. Then $w_{ijkm}=E[\xi^{(1)}_{ij}\xi^{(2)}_{km}]=E[\xi^{(1)}_{ij}]E[\xi^{(2)}_{km}]=w^{(1)}_{ij}w^{(2)}_{km}$ that is $W_{S}=W_{S_{1}}\otimes W_{S_{2}}$. We also note that $||W||=||W^{(1)}||\;||W^{(2)}||$ and hence the state of $S$ is factorized, $\omega=\omega^{(1)}\otimes\omega^{(2)}$. Here $\omega^{(1)}$ and $\omega^{(2)}$ are the states of the networks $S_{1}$ and $S_{2}$, $\omega^{(1)}=\omega_{S_{1}}$ and $\omega^{(2)}=\omega_{S_{2}}$, as they were defined in section 2.
For a state $\omega\in\Omega_{S}$, the states of its subsystems are defined as follows,
$\omega^{(1)}=(\omega^{(1)}_{ij}=\sum_{km}w_{ijkm}),\;\omega^{(2)}=(\omega^{(1)}_{km}=\sum_{ij}w_{ijkm}),$
and $\omega^{(p)}\in\Omega_{S_{p}},p=1,2$. We remark that in terms of random signaling variables, for example, $\omega^{(1)}_{ij}=E[\xi^{(1)}_{ij}(\sum_{km}\xi^{(2)}_{km})]$, and the random variable $\Xi^{(2)}=\sum_{km}\xi^{(2)}_{km}$ can be interpreted as the intensity of signaling within the network $S_{2}$. Hence, the network $S_{1}$ can be considered as open quantum-like system interacting with environment - the network $S_{2}$, and vice verse. In this model $S_{1}$ “feels” not signaling between the individual neurons of $S_{2}$, but the integral intensity of signaling.
One can proceed further by employing the general formalism of measurement instruments in tensor products.
7 Concluding remarks
As was pointed out, quantum-like modeling is actively developed as a phenomenological approach. But it really cries for coupling with brain’s functioning, with activity of the networks of communicating neurons. In this article the basics of such coupling were established within the GPT model based on the representation of networks by ordered weighted graphs that in turn are described by the matrices of weights for intensities of signaling between neurons. The graph-matrix GPT model (Model 1) has classical state space, a polygon, and its quantum-like properties are generated by the calculus of measurement instruments.
We presented some quantum-like effects, the order, interference (disjunction), and non-repeatability effects, within the graph-matrix GPT model. The compound systems, correlated neuronal networks, are considered only briefly and we plan to study quantum-like properties related to entanglement of the states of neuronal networks in one of future publications.
Acknowledgments
A.K. was partially supported by the EU-grant CA21169 (DYNALIFE), by JST, CREST Grant Number JPMJCR23P4, Japan, and visiting professor fellowship to Tokyo University of Science (April 2024). M.O. was partially supported by JSPS KAKENHI Grant Numbers JP24H01566, JP22K03424, JST CREST Grant Number JPMJCR23P4, RIKEN TRIP Initiative (RIKEN Quantum), and the Quantinuum–Chubu University Collaboration in 2023–2024.